OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line y=5/2x+2

OpenStudy (anonymous):

When two lines are perpendicular, the slope of one is the negative reciprocal of the other. so If the slope of one line is \(m\) , the slope of the other is \(-\frac{1}{m}\)

OpenStudy (anonymous):

Would the answer be -5/2

OpenStudy (anonymous):

slope of \(y=\frac{5}{2}x+2\) is \(5/2\) so slope of the line that is perpendicular to this line is \[-\frac{1}{\frac{5}{2}}=?\]

OpenStudy (anonymous):

the slope will be (-2/5).......... your original slope is (5/2) to find the perpendicular slope you take the reciprocal AND the opposite sign (either + positive or negative -)

OpenStudy (anonymous):

to check your answer simply multiply your original slope by your newly found slope and it should equal (-1) (5/2)*(-2/5)=-1

OpenStudy (anonymous):

Thank You